The domain of x is {-1, 0, 2}. Solve the equation: 6x - 0.5y = 3




Answer by josgarithmetic(39617)

6x - 0.5y = 3
12x-y=6
-y=12x+6
y=-12x-6

The given domain will make the range {6, -6, -30}.